Question 10 (1 point) For each of the following t-accounts, calculate the missing amount and enter it in the space provided. Sales 7/9 Transaction B 7/18 Transaction C 1,500 7/1 Beginning balance 2,000 7/6 Transaction A 20,000 7/28 Transaction D 3,400 28,200 Ending balance Make sure your answer does not include a $ (dollar sign) or , (comma). It should be in the format of a basic number; such as 55000For each of the following t-accounts, calculate the missing amount and enter it in the space provided Cash 5/1 Beginning balance 5/18 Transaction B ?5/16 Transaction A 5,000 5/19 Transaction C 2,000 25,000 Ending balance 65,000 Make sure your answer does not include a $ (dollar sign) or (comma). It should be in the format of a basic number; such as 55000.For each of the following t-accounts, calculate the missing amount and enter it in the space provided Unearned Revenue 3/1 Beginning balance 16,800 3/17 TransactionB ?3/29Transaction D 30,000 25,000 3,500 35,500 3/16 TransactionA 3/28 Transaction C Ending balance Make sure your answer does not include a $ (dollar sign) or (comma). It should be in the format of a basic number; such as 55000
Answer #1
1
Missing amount 8300 =28200-3400-20000+1500+2000
2
Missing amount 87000 =65000-5000+2000+25000
3
Missing amount 6200 =30000+25000+3500-16800-35500